Home‎ > ‎Crew 155‎ > ‎

1028-CommandersReport

Commander Report  
28 OCT 2015 -- Sol 5
Paul Bakken


*** BEGIN TRANSMISSION ***

Greetings Mission Support!

Crew 155 will file a Commander Report, Engineer Report, and Journalist Report today. EVAs were as described during comms last night. No separate report will be filed for the EVAs.  Instead, I will cover the pertinent details in this Commander report.

The crew woke up today at 0730 hours to the delightful sound of Reveille, followed by the original version of Space Oddity by David Bowie.  After a breakfast together and morning brief, we began the work for the day.

Today is "Ochi Day" in Greece and Cyprus.  ENG Palpanis explained to us the history of this Greek holiday.

The primary focus of the day was training the crew on EVA suit use and conducting a number of short EVAs to allow the crew to become familiar with the operation and limitations of the suits, LSU packs, and helmets.  These EVAs also gave crew a chance to try out the different combinations of equipment in order to identify the particular combinations that worked best for each individual.  This has been especially important, since the sizes of our crew members vary from 5'5" to 6'6".

The EVAs included some light work clearing the soil near the old back porch and practicing Engineering Report tasks.  We were fortunate to have some of the EVA activity documented by our Crew Journalist, Cassie Klos.

I installed the prototype airlock SCADA system under Sim conditions with the assistance of HSO Harrison and GHO Nguyen.  The system works well and adds quite a bit of verisimilitude to the Sim experience.

While not on EVA, XO Silva-Martinez and ENG Palpanis finished up the electrical systems tracing project.  We have identified a number of inoperative sockets and switches, as well as a few outlets that have been wired incorrectly.  For example, some of the outlets in the EVA prep room have the ground and neutral switched around.  I have the experience and equipment need to fix these, so I would like to request permission from Mission Support to get these outlets up to code.

XO Silva-Martinez and ENG Palpanis have also created a floor plan style diagram that shows which electrical services are on which breaker circuit.  We will get this polished and make it available as a general resource to the MDRS Engineering Team.

CJ Klos and XO Silva-Martinez organized and cleaned the lab and engineering areas on Deck 1.  CJ Klos also performed more painting touch ups on the walls of Deck 1 and Deck 2.

HSO Harrison set traps for the four legged vermin last night.  This morning, we found that the peanut butter had been eaten off of all but one of the traps without triggering them.  The remaining trap was also baitless, but had been triggered.  Contemptuously, I suspect, since the trap was unoccupied.  I am beginning to think that the MDRS has unwittingly been applying selective evolutionary pressure to the rodent population in the Hab over the last decade, and has mainly succeeded in breeding an unusually intelligent and snarky breed of mice.  Efforts at evicting these unwanted tenants continue.  But seriously, who thought it was a good idea to bring mice to Mars in the first place?!

Crew 155 finished fixing the EVA equipment that we could. There are five functional LSU packs, three helmets, and the one functional integrated model unit.  We have only four functional headsets for the radios.  The rest are garbage, and cannot be easily fixed.  If I recall correctly, this has been a recurrent challenge at the MDRS.  The type of headsets used tend not to last very long.  When I return to Earth, I will investigate options for a more durable model.

Finally, I got my prototype ozone laundering machine assembled and tested today. My crew members were intrigued by the device and look forward to participating in the laundering experiment.

XO Silva-Martinez is cooking evening meal, and it smells delicious.  Tonight, after Hab PMCS is complete, we plan to have a movie night.

Tomorrow, Crew 155 will again conduct multiple, short-duration EVAs in the vicinity of the Hab to continue the familiarization process, move more soil from the old back porch area, and possibly assemble parts of the pressurized tunnel system.

It is hard to believe that we have already been here for almost a week. Crew have settled into a good routine, and we expect to accomplish many more of the things on the "fix it" list.

That's all for now.  The crew sends best wishes to friends and family on Earth.

Respectfully submitted,

Paul Bakken    
Commander, MDRS Crew 155

*** END TRANSMISSION ***